# Fast rotations in galaxies at cosmic noon indicate central concentration of stars, dark matter or massive black holes

### v1

@@ Abstract
Gaseous and stellar discs are a predominant morphological feature in local galaxies. Due to their regular structure and kinematics, they allow us to obtain information about their mass distribution and different mass components (bosma78; rubin80; vanalbada85; mcgaugh04; delgado-serrano10; lelli16b). This provides a unique opportunity to study the interplay between baryonic and dark matter. In the local Universe, studies of neutral atomic hydrogen (HI) are the best for tracing the rotation curves of galaxies due to the radial extent of the HI emission and the quality of the interferometric data. These studies led the way to detailed mass models of rotation curves of hundreds of galaxies in the local Universe, ranging from dwarf galaxies to giant spirals (lelli16; mancerapina22a; mancerapina22b; biswas23). This is possible due to the ability of a cold gas component, such as HI, to trace the circular speed and, effectively, the gravitational potential as a function of galactocentric radius. Observational constraints on the mass distribution of the baryonic components, namely gas and stars, allow then to infer the properties of the dark matter halo (deblok10; penarrubia10; tulin18; read19; posti19; posti21; nadler23; banareshernanez23; mancera2025). Some of these studies have also opened debates over the shape of dark matter halos, such as the well known cusp-core problem, the nature of dark matter itself, and the impact of feedback on the stellar-to-halo mass ratios (tulin18; read19; mancerapina22a).
Deriving the properties of dark matter halos at 0.5 through the analysis of rotation curves is significantly more challenging than at . The first difficulty is obtaining accurate kinematics, as the HI emission is no longer available due to the relatively low sensitivity of current radio telescopes (verheijen07; jaffe12; fernandez16; chowdhury22). At , kinematic studies of galaxies have relied on alternative tracers, mostly emission lines that trace warm ionised gas, such as H, [OII] and [OIII] (wisnioski15; turner17; forster18; puglisi23; Danhaive2025). These emission lines are very bright and, therefore, they are the best tracer for obtaining large samples of hundreds of galaxies. However, the gravitational potential may not be optimally characterised, as the motions of the warm gas might be influenced by non-circular motions driven by Active Galactic Nuclei (AGN) or stellar feedback (e.g., levy18; girard21; rizzo24; kohandel24; Phillips_2025).
@@
As a consequence of the above, in previous attempts, the total gas mass has typically been fixed based on average values of the gas fraction expected for the star-forming population or estimated from scaling relations that correlate the molecular gas mass to the star formation rate (e.g., tacconi18). Additionally, the spatially-resolved gas mass profile has been assumed to follow the distribution of the warm gas (e.g., genzel13; uebler18; genzel20; puglisi23; Ciocan_2025; Sharma_2025).
A further complication arises when either the data quality (e.g., limited spatial or spectral resolution) or the intrinsic nature of the observations (e.g., slitless spectroscopy) prevents a reliable assessment of whether a galaxy hosts a disk whose gas kinematics is dominated by circular motions (degraaff24; Lee_2025; Danhaive_2025b). As a result, the inferred parameters of the dark-matter halo can become highly uncertain or even unreliable.
The above assumptions introduce several uncertainties, and an ideal way to overcome them is to combine high spectral and angular resolution observations of emission lines tracing cold gas from ALMA with rest-frame near-infrared observations from JWST. However, at , the only way to obtain spatially resolved kinematics of cold gas is to target CO or [CI] transitions. These lines are fainter than [CII] and, therefore, spatially-resolved observations of cold gas are available only for a handful of galaxies in this redshift range (e.g. noble19; molina19; kaasinen20; xiao22; lelli23).
In a recent effort to address the lack of high-quality emission line observations of cosmic noon galaxies, the ALMA Archival Large Programme to Advance Kinematic Analysis (ALMA-ALPAKA) project was presented in rizzo23. This is the most comprehensive sample of resolved cold gas observations in galaxies at and it attempts to bridge the gap in observational cold gas kinematic studies using medium to high-resolution ( 0.2") observations of CO or [CI] emission lines from the ALMA archive.
@@
The paper is structured as follows: in Section 2, we describe the data used; in Section 3, we describe the methodology employed in the analysis of the rotation curves; in Section 4, we present our main results regarding the mass decomposition of the rotation curves; in Section 5, we discuss our main findings; in Section 6, we summarise our results and their implications within the field. Throughout this work, we adopt a CDM cosmology from the 2018 Planck results (planck20).
@@ 2.1 ALMA-ALPAKA subsample
In rizzo23, the 28 galaxies in the ALPAKA sample were classified according to their kinematics, and it was found that 19 galaxies are regularly rotating discs. The remaining galaxies were classified as merging systems (2/28) and uncertain (7). The latter means that the quality of the data is not good enough to robustly classify the galaxies as a disc or an interacting system. For the present work, we selected only galaxies that: i) were classified as regular discs; ii) have at least three resolution independent elements over the major axis; iii) have public data from JWST/NIRCAM; (iv) are at . We note that, despite some galaxies having Hubble Space Telescope (HST) imaging and/or an estimate of the stellar mass through Spectral Energy Distribution (SED) fitting, it does not necessarily mean that there is a good constraint on the stellar distribution. This is because the reddest HST band covers, at most, the rest-frame optical and not the near-infrared.
Moreover, for relatively dusty galaxies, the light distribution is significantly attenuated, and we cannot robustly recover the stellar profile (polletta24). Therefore, to reduce this bias, our final subsample contains the three galaxies for which there is coverage of the rest-frame near-infrared. This emission comes from the bulk of the stellar populations and it is typically less affected by dust attenuation. In Table 1, we list the ID and redshift of the ALPAKA subsample used in this work.
@@ 2.2 ALMA data
The ALMA datacubes used in this work were calibrated and imaged as described in rizzo23. For this work, we collapsed the continuum subtracted datacubes in the velocity direction to create moment 0 maps of the CO/[CI] emission line. The velocity range of the original cubes extends for at least . These maps are centred in the galaxy and have wide enough fields of view to encompass the whole emission. We used these maps to model the gas surface brightness profile. We also used the rotation velocity profiles as a function of radius, obtained through the kinematic models with 3DBAROLO, reported in rizzo23. In Figure 1, we overlay the gas moment 0 maps on top of the rest-frame near-infrared data of our sample (discussed below).
@@ 2.3 Rest-frame optical and near-infrared data
For our three targets, we use publicly available JWST/NIRCam data from NIRCam, processed as part of the DAWN JWST Archive (DJA, Valentino_2023; Heintz_2025). To have homogeneous coverage in the analysis of the stellar continuum distribution, we select the JWST filters that cover the rest-frame emission at 1.5 – 2 m, corresponding to F277W for ID1 and F444W for ID3 and I13 (Table 1). In Figure 1, we show a colour-composited image using the HST and JWST images and in Table 1 we report the rest-frame wavelengths of the filters used.
@@ 3.1.1 Gas distribution
To fit the gas distribution of the three ALPAKA targets, we used the methodology described in romanoliveira24 assuming that they follow an exponential profile, corresponding to a Sérsic profile (sersic68) with the Sérsic index fixed at 1. In addition to the parameters defining the profile (i.e., intensity at the effective radius , effective radius ), we fitted the geometrical parameters: centre, defined by and , ellipticity , and position angle . We fitted the 2D surface brightness model using DYNESTY, a python package for estimating Bayesian posteriors and evidences with the Dynamic Nested Sampling algorithm (skilling04; skilling06; speagle20; koposov22). Before fitting the data, we convolve the model to match the same spatial resolution as the data. Moreover, similarly to romanoliveira24 and marasco19, we modified the minimisation function to better represent the errors in the data. We used a chi-square likelihood with residuals weighted by an effective uncertainty that depends on the asymmetry of the galaxy, such that: , where is the data and is the model. The effective uncertainty () is defined as the standard deviation of the residual map obtained by rotating the data map by 180∘ and subtracting it from the original, non-rotated data map. This provides a more realistic (and larger) estimate of the errors in the best-fit parameters, as the source of mismatch in the fitting is primarily due to deviation from the axisymmetry assumed in the model rather than the background noise in the data. In Table 2, we summarise the priors used in the fitting of the gas surface brightness.
@@ 3.1.2 Stellar distribution
The code used to model the gas distribution (Sect. 3.1.1) was originally developed for ALMA data, where the beam is well approximated by a Gaussian kernel. In contrast, the JWST point spread function (PSF) is considerably more complex and cannot be accurately described by a Gaussian profile. Because adapting and running the gas-distribution code with the full JWST PSF would be computationally expensive, we instead employed GALFIT (Peng_2010) to model the stellar surface brightness. To characterise the JWST PSF, we identified an unsaturated star within 30” of each target galaxy. As shown in Figure 1, for ID3 and ID13 the modelling also required simultaneous fits to nearby foreground and background sources in order to obtain reliable structural parameters for the main galaxies.
For the three targets in our sample, we fitted the main galaxy with two components (one for the bulge and another for the disc) and an additional flat component for the sky background. For ID3 and ID13, we also fit extra components for the foreground and background galaxies, for which we assumed a Sérsic profile (sersic68) and let the total intensity, effective radius and Sérsic index free to vary.
@@ 3.2 Circular Speed
where is the correction for asymmetric drift, and are the rotation velocity and velocity dispersion of the gas obtained from kinematic modelling with 3DBAROLO, and is the surface density of the gas. Here, we assume that the disc thickness is radially constant. To see this derivation more explicitly, we refer to iorio17. We note that in galaxies dominated by rotational motion (), this correction is usually minimal. However, different regions of the same galaxy may have different pressure support and will be affected differently as the correction is applied to each point in the radial profile. This can have an impact on the overall shape of the rotation curve.
@@
where is the disc-scale length of the exponential gas disc that is defined as . To apply this correction, we use and from the kinematic values derived in rizzo23, we then need to derive for each galaxy as we explain in Section 3.1.
@@ 3.3 Rotation Curve Decomposition
Stellar Bulge: The projected light distribution is assumed to follow a de Vaucouleurs profile. To describe the circular velocity contribution of the bulge (), we adopt the generalised deprojected spherical Sérsic profile of terzic05, which gives
@@
Stellar Disc: The stellar disc is described by a thin exponential disc (binney08), whose circular speed is given by
@@
Gas Disc: The gaseous disc is also modelled as a thin exponential disc (Eq. 7), but truncated at the observed CO/[CI] radius since no aperture correction is applied to molecular gas masses. Following binney08, the central surface density is
@@
Dark Matter Halo: For the dark matter distribution, we assume a spherical Navarro–Frenk–White (NFW) halo profile (navarro97). The circular speed is given by
@@
| NFW concentration, | Fixed | CMR (dutton14) | - | |
@@ 3.3.1 Modelling the rotation curves
We fitted the galaxy rotation curves, corrected for asymmetric drift, using Equation (5). The posterior distribution of the model parameters was explored using the nested-sampling algorithm DYNESTY (speagle20). The fits were evaluated by minimising the residuals through a chi-square likelihood, , where represents the uncertainty in the observed rotation curve.
In our fiducial models, we used the distribution of the gas and stellar surface density obtained by fitting the ALMA and JWST data. We kept them fixed while leaving their normalisation (i.e., , ) free to vary. We included the NFW halo with a concentration following the concentration mass relation (CMR) from dutton14, which is described as
@@
The mass of the dark matter halo () was calculated indirectly from a baryonic fraction parameter (), defined as the ratio between the baryonic mass () and the dark matter halo virial mass, = . The baryonic fraction parameter was kept free with a flat prior. The upper limit of the prior is given by the cosmological baryonic fraction of 18.7% from planck20, ensuring that there is a physically motivated minimal amount of dark matter contributing to the rotation curves within R200. The lower limit was set to a baryonic fraction of 10-7, which corresponds to a heavily dark matter dominated galaxy. This lower limit is motivated to set a wide enough prior range for the posteriors so that they are not affected by the edges of the prior, as the baryonic fraction tends to have a tail towards the lower end (see Figure 15).
@@
where is the CO/[CI] luminosity, is the luminosity line ratio and is the CO/[CI] luminosity-to-H2 mass conversion. We kept the product of as a free parameter in the fiducial models with a flat prior. In Section 4.2.2, instead we kept these parameters fixed as we attempted to constrain the dark matter component by fixing the baryonic components. In this case, we chose an value of 3.6 and 0.8 M for ID1 and ID3, as these are the typical values used for main-sequence and starburst galaxies, respectively (daddi10a; kirkpatrick19). For ID13, we used the [CI] (3P2 - 3P1) emission to derive the total molecular mass after assuming an factor. Using astrochemical simulations, bisbas21 suggested that the values range from 1 to 9.7 M, with the most common value being 7 M. Other observational works on both local and high- galaxies have found values that roughly agree with this range (valentino18; crocker19; valentino20b; heintz20; lelli23). Another study focused on submillimitre galaxies has found somewhat higher values of M dunne22. In Section 4.2.2, we chose a conservative of 1 M, the lower end of the range of values reported in the literature. This is because larger values of , combined with the estimated stellar mass (Table 1) would overpredict the circular speed in the outer parts of the disc.
@@
Regarding the luminosity line ratios in Section 4.2.2, we used the results in boogaard20 for the CO emission, taking the values from their Table 3 for the subsamples at for ID1 and ID3. We note that this is the best collection of line ratios for galaxies in this redshift range, but are uncertain. It contains only 12 galaxies to calibrate the line ratio models, and only three of those were observed in CO (5-4), the same transition as ID3. For ID13, as discussed in valentino18, the simple structure of neutral atomic carbon emission allows for the breaking of the temperature-density degeneracy that is not possible in the case of CO. Therefore, the excitation temperature of the gas can be directly related to the luminosity line ratio by T (schneider03; weiss03). Since we only have one [CI] transition available, we assumed a gas excitation temperature of , as found on average by other works at a similar redshift (walter11; valentino20b; henriquezbrocal22). We obtained an estimate of the L’/L’ line ratio of 0.57.
@@
For the stellar distribution, we adopted the structural parameters of the bulge and disc derived with GALFIT, keeping them fixed during the modelling. The bulge-to-disc luminosity ratio was also taken from the GALFIT decomposition. We assumed that the bulge mass-to-light ratio () is 1.4 times higher than that of the disc, following the approach of lelli16b for spiral galaxies at . As discussed in Section 5, the different mass-to-light ratios for the bulge and the disc components can be attributed to a difference in their stellar populations (e.g., schombert14) or a gradient in dust attenuation. In our fiducial fit (Section 4.2.1), we allowed the total stellar mass to vary independently of the prior stellar masses derived from SED fitting; however, we kept the relative contributions of mass in the disc and bulge fixed.

### v2

@@ Abstract
Gaseous and stellar discs are a predominant morphological feature in local galaxies. Due to their regular structure and kinematics, they allow us to obtain information about their mass distribution and different mass components (Bosma, 1978; Rubin et al., 1980; van Albada et al., 1985; McGaugh, 2004; Delgado-Serrano et al., 2010; Lelli et al., 2016a). This provides a unique opportunity to study the interplay between baryonic and dark matter. In the local Universe, studies of neutral atomic hydrogen (HI) are the best for tracing the rotation curves of galaxies due to the radial extent of the HI emission and the quality of the interferometric data. These studies led the way to detailed mass models of rotation curves of hundreds of galaxies in the local Universe, ranging from dwarf galaxies to giant spirals (Lelli et al., 2016b; Mancera Piña et al., 2022b, a; Biswas et al., 2023). This is possible due to the ability of a cold gas component, such as HI, to trace the circular speed and, effectively, the gravitational potential as a function of galactocentric radius. Observational constraints on the mass distribution of the baryonic components, namely gas and stars, allow then to infer the properties of the dark matter halo (de Blok, 2010; Peñarrubia et al., 2010; Tulin and Yu, 2018b; Read et al., 2019; Posti et al., 2019; Posti and Fall, 2021; Nadler et al., 2023; Bañares-Hernández et al., 2023; Mancera Piña et al., 2025). Some of these studies have also opened debates over the shape of dark matter halos, such as the well known cusp-core problem, the nature of dark matter itself, and the impact of feedback on the stellar-to-halo mass ratios (Tulin and Yu, 2018b; Read et al., 2019; Mancera Piña et al., 2022b).
Deriving the properties of dark matter halos at 0.5 through the analysis of rotation curves is significantly more challenging than at . The first difficulty is obtaining accurate kinematics, as the HI emission is no longer available due to the relatively low sensitivity of current radio telescopes (Verheijen et al., 2007; Jaffé et al., 2012; Fernández et al., 2016; Chowdhury et al., 2022). At , kinematic studies of galaxies have relied on alternative tracers, mostly emission lines that trace warm ionised gas, such as H, [OII] and [OIII] (Wisnioski et al., 2015; Turner et al., 2017; Förster Schreiber et al., 2018; Puglisi et al., 2023; Danhaive et al., 2025b). These emission lines are very bright and, therefore, they are the best tracer for obtaining large samples of hundreds of galaxies. However, the gravitational potential may not be optimally characterised, as the motions of the warm gas might be influenced by non-circular motions driven by Active Galactic Nuclei (AGN) or stellar feedback (e.g., Levy et al., 2018; Girard et al., 2021; Rizzo et al., 2024; Kohandel et al., 2024; Phillips et al., 2025).
@@
As a consequence of the above, in previous attempts, the total gas mass has typically been fixed based on average values of the gas fraction expected for the star-forming population or estimated from scaling relations that correlate the molecular gas mass to the star formation rate (e.g., Tacconi et al., 2018). Additionally, the spatially-resolved gas mass profile has been assumed to follow the distribution of the warm gas (e.g., Genzel et al., 2013; Übler et al., 2018; Genzel et al., 2020; Puglisi et al., 2023; Ciocan et al., 2025; Sharma et al., 2025).
A further complication arises when either the data quality (e.g., limited spatial or spectral resolution) or the intrinsic nature of the observations (e.g., slitless spectroscopy) prevents a reliable assessment of whether a galaxy hosts a disk whose gas kinematics is dominated by circular motions (de Graaff et al., 2024; Lee et al., 2025; Danhaive et al., 2025a). As a result, the inferred parameters of the dark-matter halo can become highly uncertain or even unreliable.
The above assumptions introduce several uncertainties, and an ideal way to overcome them is to combine high spectral and angular resolution observations of emission lines tracing cold gas from ALMA with rest-frame near-infrared observations from JWST. However, at , the only way to obtain spatially resolved kinematics of cold gas is to target CO or [CI] transitions. These lines are fainter than [CII] and, therefore, spatially-resolved observations of cold gas are available only for a handful of galaxies in this redshift range (e.g. Noble et al., 2019; Molina et al., 2019; Kaasinen et al., 2020; Xiao et al., 2022; Lelli et al., 2023).
In a recent effort to address the lack of high-quality emission line observations of cosmic noon galaxies, the ALMA Archival Large Programme to Advance Kinematic Analysis (ALMA-ALPAKA) project was presented in Rizzo et al. (2023). This is the most comprehensive sample of resolved cold gas observations in galaxies at and it attempts to bridge the gap in observational cold gas kinematic studies using medium to high-resolution ( 0.2") observations of CO or [CI] emission lines from the ALMA archive.
@@
The paper is structured as follows: in Section 2, we describe the data used; in Section 3, we describe the methodology employed in the analysis of the rotation curves; in Section 4, we present our main results regarding the mass decomposition of the rotation curves; in Section 5, we discuss our main findings; in Section 6, we summarise our results and their implications within the field. Throughout this work, we adopt a CDM cosmology from the 2018 Planck results (Planck Collaboration et al., 2020).
@@ 2.1 ALMA-ALPAKA subsample
In Rizzo et al. (2023), the 28 galaxies in the ALPAKA sample were classified according to their kinematics, and it was found that 19 galaxies are regularly rotating discs. The remaining galaxies were classified as merging systems (2/28) and uncertain (7). The latter means that the quality of the data is not good enough to robustly classify the galaxies as a disc or an interacting system. For the present work, we selected only galaxies that: i) were classified as regular discs; ii) have at least three resolution independent elements over the major axis; iii) have public data from JWST/NIRCAM; (iv) are at . We note that, despite some galaxies having Hubble Space Telescope (HST) imaging and/or an estimate of the stellar mass through Spectral Energy Distribution (SED) fitting, it does not necessarily mean that there is a good constraint on the stellar distribution. This is because the reddest HST band covers, at most, the rest-frame optical and not the near-infrared.
Moreover, for relatively dusty galaxies, the light distribution is significantly attenuated, and we cannot robustly recover the stellar profile (Polletta et al., 2024). Therefore, to reduce this bias, our final subsample contains the three galaxies for which there is coverage of the rest-frame near-infrared. This emission comes from the bulk of the stellar populations and it is typically less affected by dust attenuation. In Table 1, we list the ID and redshift of the ALPAKA subsample used in this work.
@@ 2.2 ALMA data
The ALMA datacubes used in this work were calibrated and imaged as described in Rizzo et al. (2023). For this work, we collapsed the continuum subtracted datacubes in the velocity direction to create moment 0 maps of the CO/[CI] emission line. The velocity range of the original cubes extends for at least . These maps are centred in the galaxy and have wide enough fields of view to encompass the whole emission. We used these maps to model the gas surface brightness profile. We also used the rotation velocity profiles as a function of radius, obtained through the kinematic models with 3DBAROLO, reported in Rizzo et al. (2023). In Figure 1, we overlay the gas moment 0 maps on top of the rest-frame near-infrared data of our sample (discussed below).
@@ 2.3 Rest-frame optical and near-infrared data
For our three targets, we use publicly available JWST/NIRCam data from NIRCam, processed as part of the DAWN JWST Archive (DJA, Valentino et al., 2023; Heintz et al., 2025). To have homogeneous coverage in the analysis of the stellar continuum distribution, we select the JWST filters that cover the rest-frame emission at 1.5 – 2 m, corresponding to F277W for ID1 and F444W for ID3 and I13 (Table 1). In Figure 1, we show a colour-composited image using the HST and JWST images and in Table 1 we report the rest-frame wavelengths of the filters used.
@@ 3.1.1 Gas distribution
To fit the gas distribution of the three ALPAKA targets, we used the methodology described in Roman-Oliveira et al. (2024) assuming that they follow an exponential profile, corresponding to a Sérsic profile (Sersic, 1968) with the Sérsic index fixed at 1. In addition to the parameters defining the profile (i.e., intensity at the effective radius , effective radius ), we fitted the geometrical parameters: centre, defined by and , ellipticity , and position angle . We fitted the 2D surface brightness model using DYNESTY, a python package for estimating Bayesian posteriors and evidences with the Dynamic Nested Sampling algorithm (Skilling, 2004, 2006; Speagle, 2020; Koposov et al., 2022). Before fitting the data, we convolve the model to match the same spatial resolution as the data. Moreover, similarly to Roman-Oliveira et al. (2024) and Marasco et al. (2019), we modified the minimisation function to better represent the errors in the data. We used a chi-square likelihood with residuals weighted by an effective uncertainty that depends on the asymmetry of the galaxy, such that: , where is the data and is the model. The effective uncertainty () is defined as the standard deviation of the residual map obtained by rotating the data map by 180∘ and subtracting it from the original, non-rotated data map. This provides a more realistic (and larger) estimate of the errors in the best-fit parameters, as the source of mismatch in the fitting is primarily due to deviation from the axisymmetry assumed in the model rather than the background noise in the data. In Table 2, we summarise the priors used in the fitting of the gas surface brightness.
@@ 3.1.2 Stellar distribution
The code used to model the gas distribution (Sect. 3.1.1) was originally developed for ALMA data, where the beam is well approximated by a Gaussian kernel. In contrast, the JWST point spread function (PSF) is considerably more complex and cannot be accurately described by a Gaussian profile. Because adapting and running the gas-distribution code with the full JWST PSF would be computationally expensive, we instead employed GALFIT (Peng et al., 2010) to model the stellar surface brightness. To characterise the JWST PSF, we identified an unsaturated star within 30” of each target galaxy. As shown in Figure 1, for ID3 and ID13 the modelling also required simultaneous fits to nearby foreground and background sources in order to obtain reliable structural parameters for the main galaxies.
For the three targets in our sample, we fitted the main galaxy with two components (one for the bulge and another for the disc) and an additional flat component for the sky background. For ID3 and ID13, we also fit extra components for the foreground and background galaxies, for which we assumed a Sérsic profile (Sersic, 1968) and let the total intensity, effective radius and Sérsic index free to vary.
@@ 3.2 Circular Speed
where is the correction for asymmetric drift, and are the rotation velocity and velocity dispersion of the gas obtained from kinematic modelling with 3DBAROLO, and is the surface density of the gas. Here, we assume that the disc thickness is radially constant. To see this derivation more explicitly, we refer to Iorio et al. (2017). We note that in galaxies dominated by rotational motion (), this correction is usually minimal. However, different regions of the same galaxy may have different pressure support and will be affected differently as the correction is applied to each point in the radial profile. This can have an impact on the overall shape of the rotation curve.
@@
where is the disc-scale length of the exponential gas disc that is defined as . To apply this correction, we use and from the kinematic values derived in Rizzo et al. (2023), we then need to derive for each galaxy as we explain in Section 3.1.
@@ 3.3 Rotation Curve Decomposition
Stellar Bulge: The projected light distribution is assumed to follow a de Vaucouleurs profile. To describe the circular velocity contribution of the bulge (), we adopt the generalised deprojected spherical Sérsic profile of Terzić and Graham (2005), which gives
@@
Stellar Disc: The stellar disc is described by a thin exponential disc (Binney and Tremaine, 2008), whose circular speed is given by
@@
Gas Disc: The gaseous disc is also modelled as a thin exponential disc (Eq. 7), but truncated at the observed CO/[CI] radius since no aperture correction is applied to molecular gas masses. Following Binney and Tremaine (2008), the central surface density is
@@
Dark Matter Halo: For the dark matter distribution, we assume a spherical Navarro–Frenk–White (NFW) halo profile (Navarro et al., 1997). The circular speed is given by
@@
| NFW concentration, | Fixed | CMR (Dutton and Macciò, 2014) | - | |
@@ 3.3.1 Modelling the rotation curves
We fitted the galaxy rotation curves, corrected for asymmetric drift, using Equation (5). The posterior distribution of the model parameters was explored using the nested-sampling algorithm DYNESTY (Speagle, 2020). The fits were evaluated by minimising the residuals through a chi-square likelihood, , where represents the uncertainty in the observed rotation curve.
In our fiducial models, we used the distribution of the gas and stellar surface density obtained by fitting the ALMA and JWST data. We kept them fixed while leaving their normalisation (i.e., , ) free to vary. We included the NFW halo with a concentration following the concentration mass relation (CMR) from Dutton and Macciò (2014), which is described as
@@
The mass of the dark matter halo () was calculated indirectly from a baryonic fraction parameter (), defined as the ratio between the baryonic mass () and the dark matter halo virial mass, = . The baryonic fraction parameter was kept free with a flat prior. The upper limit of the prior is given by the cosmological baryonic fraction of 18.7% from Planck Collaboration et al. (2020), ensuring that there is a physically motivated minimal amount of dark matter contributing to the rotation curves within R200. The lower limit was set to a baryonic fraction of 10-7, which corresponds to a heavily dark matter dominated galaxy. This lower limit is motivated to set a wide enough prior range for the posteriors so that they are not affected by the edges of the prior, as the baryonic fraction tends to have a tail towards the lower end (see Figure 15).
@@
where is the CO/[CI] luminosity, is the luminosity line ratio and is the CO/[CI] luminosity-to-H2 mass conversion. We kept the product of as a free parameter in the fiducial models with a flat prior. In Section 4.2.2, instead we kept these parameters fixed as we attempted to constrain the dark matter component by fixing the baryonic components. In this case, we chose an value of 3.6 and 0.8 M for ID1 and ID3, as these are the typical values used for main-sequence and starburst galaxies, respectively (Daddi et al., 2010; Kirkpatrick et al., 2019). For ID13, we used the [CI] (3P2 - 3P1) emission to derive the total molecular mass after assuming an factor. Using astrochemical simulations, Bisbas et al. (2021) suggested that the values range from 1 to 9.7 M, with the most common value being 7 M. Other observational works on both local and high- galaxies have found values that roughly agree with this range (Valentino et al., 2018; Crocker et al., 2019; Valentino et al., 2020; Heintz and Watson, 2020; Lelli et al., 2023). Another study focused on submillimitre galaxies has found somewhat higher values of M Dunne et al. (2022). In Section 4.2.2, we chose a conservative of 1 M, the lower end of the range of values reported in the literature. This is because larger values of , combined with the estimated stellar mass (Table 1) would overpredict the circular speed in the outer parts of the disc.
@@
Regarding the luminosity line ratios in Section 4.2.2, we used the results in Boogaard et al. (2020) for the CO emission, taking the values from their Table 3 for the subsamples at for ID1 and ID3. We note that this is the best collection of line ratios for galaxies in this redshift range, but are uncertain. It contains only 12 galaxies to calibrate the line ratio models, and only three of those were observed in CO (5-4), the same transition as ID3. For ID13, as discussed in Valentino et al. (2018), the simple structure of neutral atomic carbon emission allows for the breaking of the temperature-density degeneracy that is not possible in the case of CO. Therefore, the excitation temperature of the gas can be directly related to the luminosity line ratio by T (Schneider et al., 2003; Weiß et al., 2003). Since we only have one [CI] transition available, we assumed a gas excitation temperature of , as found on average by other works at a similar redshift (Walter et al., 2011; Valentino et al., 2020; Henríquez-Brocal et al., 2022). We obtained an estimate of the L’/L’ line ratio of 0.57.
@@
For the stellar distribution, we adopted the structural parameters of the bulge and disc derived with GALFIT, keeping them fixed during the modelling. The bulge-to-disc luminosity ratio was also taken from the GALFIT decomposition. We assumed that the bulge mass-to-light ratio () is 1.4 times higher than that of the disc, following the approach of Lelli et al. (2016a) for spiral galaxies at . As discussed in Section 5, the different mass-to-light ratios for the bulge and the disc components can be attributed to a difference in their stellar populations (e.g., Schombert and McGaugh, 2014) or a gradient in dust attenuation. In our fiducial fit (Section 4.2.1), we allowed the total stellar mass to vary independently of the prior stellar masses derived from SED fitting; however, we kept the relative contributions of mass in the disc and bulge fixed.